AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME, Page 25 of 33  ELOG logo
New entries since:Thu Jan 1 01:00:00 1970
ID Date Author Subject Text Attachments
  179   Thu Jan 27 15:24:02 2022 OH, JMAccess to Motors PI1. ssh <user>@lxpool.gsi.de
2. ssh <user or carme>@atppc022.gsi.de
3. ssh pi@CARMEmotorspi (Standard npg password)
  
  177   Thu Jan 27 15:19:18 2022 OH, JMEnabling internet access on the carme-gsi serverTo use both anydesk and firefox on the carme
server require an internet connection.
  
  176   Thu Jan 27 14:04:34 2022 OH, JMRemote access to CARME server1. ssh into lxpool network: ssh <user>@lxpool.gsi.de
2. ssh into atppc022.gsi.de: ssh carme@atppc022.gsi.de
(This is an rpi access server. Don't run
  
  175   Wed Jan 26 19:16:18 2022 JM, OHWater cooling, FEE MAC addressWater cooling has been connected to the
FEE's. Flow meter installed to outlet
manifold by way of two right angled pieces.
 20220126_162435.jpg20220126_162448.jpg20220126-142809.jpg 
  174   Sat Jan 15 15:38:25 2022 JM, OH, TDFEE installation, cable managementAll 8 FEEs have been installed in their
holders.

The damaged feedthrough does not
 20220114_104819.jpg20220114_105152.jpg20220114_105213.jpg20220115_133808.jpg20220115_133820.jpg 
  173   Sat Jan 15 11:28:10 2022 JMInventory 


Vaccum
DN250CF 1x gaskets
    
DN200CF 5x gaskets  
DN160CF 1x gaskets,
 20220115_162649.jpg20220115_162712.jpg 
  172   Fri Jan 14 10:00:18 2022 TD, OH, JMFriday 14 January
11.00 Per https://elog.ph.ed.ac.uk/CARME/170
and https://elog.ph.ed.ac.uk/CARME/171 issues
  
  171   Thu Jan 13 10:33:24 2022 JM, OH, TDAdaptor Board misallignmentThe ERNI connector is not central on the
adaptor boards causing misallignment with
the FEE. 
 20220112_160951.jpg20220113_145105.jpg20220113_145147.jpg 
  170   Wed Jan 12 20:07:20 2022 JM, OH, TDFEE holder installation4 FEE holders installed on CARME. Mounting
procedure devised due to limited working
space to install.
 10x 
  168   Tue Jan 11 12:45:15 2022 TD, JM, OHCoolant manifold installation + detector pin tests2x cooling water manifolds with 2x 8 ports
have been installed - see attachments 1 &
2
 20220111_113628.jpg20220111_113613.jpg20220111_111428.jpg 
  167   Mon Jan 10 20:52:53 2022 JM, TD, OHBaking tent removed, VI testsBaking tent, frame and trays fully
removed from CARME chamber. Removal of baking
frame required removing air lines from upper
 20220110_175547.jpg20220110_175811.jpgvi-curve.png 
  166   Tue Dec 7 09:39:59 2021 CB (remotely)Gate valves closedDry run complete.

YR09 section valves closed. Section
isolated. Manual valve between electron target
  
  165   Tue Nov 30 10:49:43 2021 CB (remotely) GVGate valve openedGSI vacuum division IE514 gauge connected.

Gleb reports gate downstream opened.
Vacuum goes up and recovers as expected.
 vacplot24.png 
  164   Sun Nov 28 15:57:45 2021 CB (remotely)Nov 2021 temperaturesAttached  211128164424_temperatures.zip 
  163   Fri Nov 26 10:01:45 2021 CB (remotely) MLPower cut in the CRYRING caveML reports power cut in CRYRING cave. Reasons
unknown. No effect on vacuum, but top valve
closed for safety reason. Ramping down temperature.
  
  162   Wed Nov 24 12:08:08 2021 CB + ML OGIon pumps onFound SIP power unit for D2000 #4 not functioning
correctly. Max current regardless of ion
element connected to.
 vacplot24.pngvacplot.pngCARMEtemps.png 
  161   Tue Nov 23 08:49:33 2021 CB + MLNEG activation completedFound temperatures and pressure stabilised
overnight. See attached.

12:00 Began activation of 3x UHV1400
 vacplot24.pngCARMEtemps.pngvacplot24-1.pngCARMEtemps-1.png 
  160   Mon Nov 22 08:06:01 2021 RSS CB + MLNEG activation starts08:20 Temperature of channels 1-3 and 5-6
were reduced to 120 degrees. Immediately
after, the conditioning of the NEGs was stopped
 IMG_5674.JPGCARMEtemps.pngvacplot.png 
  159   Fri Nov 19 12:03:51 2021 CB RSSConditionings NEGsIncreased conditioning temperature in D2000
#3 and 4 (behind detectors) to 150C from
130/140C. No appreciable change in temperature.
 CARMEtemps.pngvacplot.png 
  158   Thu Nov 18 13:50:51 2021 CB RSSThermocamera imagesUsed thermocamera provided by GSI beam
diagnostic division to take some pictures
of YR09. See attached.
 14x 
ELOG V3.1.4-unknown